OpenStudy (anonymous):

0< a < c < 6 < g < k What is the least possible value of a + k? These integer problems always get me . I have no idea how to approach it . Please explain

OpenStudy (anonymous):

We want the smallest a and k we can find 0<a and a is an integer, the smallest a=1; since 6<g and g is an integer, the smallest g is g=7; since g=7<k and k is an integer, the smallest k=8 thus the least a+k=1+8=9